# Phase-Field simulations of dynamic wetting by the Volume-Of-Fluid method

### v1

@@ 5.2 Spreading drop
Figure 5: Comparison of interface shapes at t = 0.1 for PF simulations (black lines) and VOF toy model (red lines). (a)-(c): Effect of the variation of the geometrical factor $a$ for a fixed value of $b$. (d)-(e) Effect of the variation of $b$ for a fixed value of $a$.
The initial radius of the drop is $R_{0}=0.5$ with an initial contact angle $\theta_{0}=90^{\degree}$. The equilibrium angle is set to $\theta_{\mathrm{e}}=70^{\degree}$. The dimensionless parameters governing the flow are $\operatorname{Ca}=0.0212$, $\operatorname{Re}=3.978$, $\operatorname{Cn}=0.01$ and $\operatorname{Pe}=1$. Figure 5 shows the influence of varying the parameters $a$ and $b$ for this transient case. Each panel shows a close up of the left contact line region, with PF results for three different times, and the corresponding VoF results, using a range of values for a and b. The best match is found for $a=3$ and $b=0.2170$ confirming the values chosen in the previous case.
Figure 6: Dimensionless diameter as a function of time for the PF and VOF simulations for th best values of $a$ and $b$.
@@
Figure 7: PF and VOF interfaces at times t = 0, 0.05, 0.1, 0.15, 0.2 and 0.4. The dotted black line corresponds to the location of the computational boundary in VOF toy model.
Figure 8: Zoom around the contact line on the PF and VOF interfaces at times t = 0, 0.05, 0.1, 0.15, 0.2 and 0.4. The dotted black line corresponds to the location of the computational boundary in VOF toy model.
@@
6 Conclusion
The phase field model for dynamic wetting simulations has the advantage that it can be explicitly connected to a particular physical situation, i.e. that of a mixture of two near immiscible fluids, and thus in principle only depends on physically meaningful parameters. Computationally it is however quite expensive, and it is not suited for large scale CFD applications. On the other hand, the Volume of Fluid method which is the method of choice for such simulations, depends on ad hoc or numerical parameters for contact line motion.
@@ 6 Conclusion
7 Declaration of Interests
The authors report no conflict of interest.

### v2

@@ V.2 Spreading drop
Figure 5: Comparison of interface shapes at t = 0.1 for phase-field simulations (black lines) and volume-of-fluid toy model (red lines). (a)-(c): Effect of the variation of the geometrical factor $a$ for a fixed value of $b$. (d)-(e) Effect of the variation of $b$ for a fixed value of $a$.
The initial radius of the drop is $R_{0}=0.5$ with an initial contact angle $\theta_{0}=90\degree$. The equilibrium angle is set to $\theta_{\mathrm{e}}=70\degree$. The dimensionless parameters governing the flow are $\operatorname{Ca}=0.0212$, $\operatorname{Re}=3.978$, $\operatorname{Cn}=0.01$ and $\operatorname{Pe}=1$. Figure 5 shows the influence of varying the parameters $a$ and $b$ for this transient case. Each panel shows a close up of the left contact line region, with PF results for three different times, and the corresponding VOF results, using a range of values for a and b. The best match is found for $a=3$ and $b=0.2170$ confirming the values chosen in the previous case.
Figure 6: Dimensionless diameter as a function of time for the phase-field and toy model volume-of-fluid simulations for the best values of $a$ and $b$.
@@
Figure 7: Phase-field and toy model volume-of-fluid interfaces at times t = 0, 0.05, 0.1, 0.15, 0.2 and 0.4. The dotted black line corresponds to the location of the computational boundary in the toy model.
Figure 8: Zoom around the contact line on the phase-field and toy model volume-of-fluid interfaces at times t = 0, 0.05, 0.1, 0.15, 0.2 and 0.4. The dotted black line corresponds to the location of the computational boundary in the toy model.
@@
VI Conclusion
The phase field model for dynamic wetting simulations has the advantage that it can be explicitly connected to a particular physical situation, i.e. that of a mixture of two near immiscible fluids, and thus in principle only depends on physically meaningful parameters. Computationally it is however quite expensive, and it is not suited for large scale CFD applications. On the other hand, the volume-of-fluid method which is the method of choice for such simulations, depends on ad hoc or numerical parameters for contact line motion.